### v1

# Enhanced entanglement from quantum ergodicity

Abstract
The quantum chaos conjecture associates the spectral statistics of a quantum system with abstract notions of quantum ergodicity. Such associations are taken to be of fundamental and sometimes defining importance for quantum chaos, but their practical relevance has been challenged by theoretical and experimental developments. Here, in counterpoint, we show that ergodic dynamics can be directly utilized for the preparation of quantum states with parametrically higher entanglement than generated by maximally scrambling circuits such as random unitaries. Our setting involves quantum systems coupled via a “non-demolition” interaction of conserved charges. We derive an exact relation between the evolving entanglement of an initial product state and a measure of spectral statistics of the interacting charges in this state. This connection is explained via a notion of Krylov vector ergodicity, tied to the ability of quantum dynamics to generate orthonormal states over time. We consider exploiting this phenomenon for the preparation of approximate Einstein-Podolsky-Rosen (EPR) states between complex systems, a crucial resource for tasks such as quantum teleportation. We quantitatively show that the transfer of operators between entangled systems, which underlies the utility of the EPR state, can be performed with parametrically larger capacity for entanglement generated via ergodic dynamics than with maximal scrambling. Our analysis suggests a direct application of “ergodic” spectral statistics as a potential resource for quantum information tasks.
Introduction— The distribution of energy levels is widely believed to be of fundamental importance in characterizing complex quantum systems. The corresponding study of “spectral statistics” has evolved two primary labels to qualitatively classify quantum dynamics. If the energy levels are arranged relatively regularly, with small spectral fluctuations around a regular integer lattice, then the label “quantum chaotic” is considered appropriate for the system [1]. Typically, such near-regular configurations are seen in the eigenvalue distributions of random matrices [1, 2], as codified in the quantum chaos conjecture [3, 4, 5, 6], but also trivially in simple ergodic systems such as the harmonic oscillator. In contrast, non-“quantum chaotic” systems tend to show large spectral fluctuations, such as for a Poisson (random) distribution of points [7]. This classification is often seen to be consistent with the qualitative behavior of a wide variety of systems that are of interest in various fields, ranging from many-body physics and statistical mechanics to quantum information and quantum gravity [8, 9, 10, 11, 12, 13, 14]. Despite the phenomenological success of these heuristic labels, a fundamental open problem is to connect spectral statistics to concrete physical processes — in other words, to identify an operational answer to “why study spectral statistics?”
Conventionally, it is believed that random matrix spectral statistics is responsible for the validity of quantum statistical mechanics, via the eigenstate thermalization hypothesis [15, 16, 17, 18, 19, 20, 21] (which amounts to an assumption on the statistics of energy eigenstates rather than eigenvalues). This motivation [20], though not quite firmly established, particularly underlies the study of spectral statistics in the aforementioned fields in the form of diagnosing “many-body quantum chaos” [20, 8, 9, 10, 11, 12, 13, 14]. However, we believe that this traditional viewpoint is (by now) significantly challenged by theoretical and experimental developments. On the experimental side, recent efforts [22, 23, 24, 25] affirm that spectral statistics is most accessible in smaller systems of $\sim 10$ particles, rather than the thermodynamically large systems of e.g. $\sim 10^{23}$ particles relevant for statistical mechanics; the potential for improved thermodynamic scaling appears to be bleak. On the theoretical side, it turns out to be possible to formulate quantum statistical mechanics in a form with equivalent and even stronger implications than eigenstate thermalization, but without making any explicit reference to the energy levels or eigenstates [26] (building on mathematical approaches in semiclassical chaos [27, 28, 29, 30, 31, 32, 33, 34, 35]).
To us, it therefore appears fairly compelling that random matrix spectral statistics is not the direct mechanism behind statistical mechanics, though they often co-occur due to the complexity of the associated quantum systems (which should not be taken as evidence of a causal link, as supported by explicit counterexamples [16, 36]). Nevertheless, spectral statistics remains a widely available resource that can potentially be harnessed: generic complex quantum systems have nontrivial spectral statistics resembling those of random matrices [1, 2]. This compels us to explore if this ubiquitous resource can be put to work for some specific task in systems of, say, a few $10$s of particles as can be manipulated in present-day quantum experiments, rather than the more macroscopic considerations of statistical mechanics for which it may well remain inaccessible.
In this work, we identify a first candidate for such a task: the preparation of highly entangled states, for use in quantum information transfer between two complex systems. The relevance of spectral statistics for this task emerges from its connection to quantum dynamical ergodicity [37, 38]: the ability of quantum dynamics to generate orthonormal states over time, given sufficiently small spectral fluctuations. We show that by engineering a non-demolition coupling [39, 22] between a control system and an ergodic system, we can prepare maximally entangled Einstein-Podolsky-Rosen (EPR) states [40] with parametrically smaller error [Eq. (15)] than the generic entangling dynamics of scrambling systems. We highlight that our analysis requires us to logically separate the notions of ergodicity and scrambling (sometimes loosely associated with each other [12, 8]), to respectively refer to orthonormal state generation in the Hilbert space [37, 38], and strongly entangling dynamics [41, 42, 43, 44, 45]. In our case, the enhanced entanglement from ergodicity is accompanied by a significantly increased capacity of the entangled state to perform tasks such as quantum teleportation [46], compared to states generated by a typical scrambler [Eq. (22)]. The details of this approach are described below.
Information transfer— Let us first review the task at hand. For two systems with Hilbert spaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ each of dimension $d$, a maximally entangled EPR state [40, 46, 47] must be of the form:
$\lvert\mathrm{EPR}\rangle=\frac{1}{\sqrt{d}}\sum_{k=0}^{d-1}\lvert k\rangle_{1}\otimes\lvert k^{\prime}\rangle_{2},$ (1)
where $\lvert k\rangle_{1}$ is an orthonormal basis for $\mathcal{H}_{1}$, as is $\lvert k^{\prime}\rangle_{2}$ for $\mathcal{H}_{2}$. After preparing this state, one can move $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ to physically very different locations, and still transfer quantum information between the two locations via local operations and classical communication [47], for example in quantum teleportation [46, 47, 48, 49]. This is due to the following crucial property of the EPR state, mathematically allowing the transfer of any operator $O_{1}$ acting on $\mathcal{H}_{1}$ to an equivalent operator $O_{2}$ acting on $\mathcal{H}_{2}$:
$O_{1}\lvert\mathrm{EPR}\rangle=O_{2}^{T}\lvert\mathrm{EPR}\rangle.$ (2)
Here, T denotes a transpose of matrix elements, with $O_{2}$ being defined by the requirement that $O_{1}$ and $O_{2}$ have identical matrix elements in the respective $k$ and $k^{\prime}$ bases: $\langle j\rvert O_{1}\lvert k\rangle_{1}=\langle j^{\prime}\rvert O_{2}\lvert k^{\prime}\rangle_{2}=\langle k^{\prime}\rvert O_{2}^{T}\lvert j^{\prime}\rangle_{2}$.
It is theoretically straightforward to prepare an EPR state between (say) many qubits using Hadamard and CNOT gates [47]. However, a lack of sufficient control in complex experimental platforms necessitates more limited system-specific protocols [48, 49, 50], in which generating maximal entanglement is often challenging [51, 52, 53]. A more generic alternative is to use some sufficiently complex “scrambling” interaction [54] between a smaller system $\mathcal{H}_{1}$ and a larger system, whichever is available in a given platform, to generate highly entangled states. Any such entangled state can be written [55, 56, 47] in the following form, where $\mathcal{H}_{2}$ is now some appropriate subspace of the larger system:
$\lvert\Psi\rangle=\frac{1}{\sqrt{d}}\sum_{k=0}^{d-1}\lvert k\rangle_{1}\otimes\lvert\psi_{k}\rangle_{2}.$ (3)
Here, $\lvert k\rangle_{1}$ remains an orthonormal basis for $\mathcal{H}_{1}$, but the $\lvert\psi_{k}\rangle_{2}$ are no longer necessarily orthonormal. Clearly, the closer the $\lvert\psi_{k}\rangle_{2}$ are to forming an orthonormal basis for $\mathcal{H}_{2}$, the more accurately the highly entangled state can emulate the information transfer property (2) of an EPR state.
From this viewpoint, our main result is that ergodic dynamics, where it can be harnessed, can generate entangled states (3) in which the $\lvert\psi_{k}\rangle_{2}$ are much closer to being an orthonormal basis for $\mathcal{H}_{2}$ than generic scrambling dynamics in many-body systems. In particular, for specific “highly ergodic” initial states, our protocol generates demonstrably more entanglement than even infinite temperature “maximal” scramblers such as random quantum circuits; for more generic initial states, the advantage in entanglement is limited to being over finite temperature scramblers such as many complex local Hamiltonians.
Krylov vector ergodicity— Our setting involves a system [47] Alice with a Hilbert space $\mathcal{H}_{A}$ of dimension $d_{A}=d$ (replacing $\mathcal{H}_{1}$ above), which contains the information to be encoded. Another system Bob with a $(d_{B}\geq d_{A})$-dimensional Hilbert space $\mathcal{H}_{B}$, is one to which we would like to transfer this information (in the subspace $\mathcal{H}_{2}$). By assumption, Alice’s system is simple and allows the application of all desired quantum operations, such as (but not necessarily) a quantum computer. Bob’s system is complex, and the best we can do is apply a certain Hamiltonian $H_{B}$. Specifically, we couple the two systems by a non-demolition interaction (see App. A for generalizations):
$H=N_{A}\otimes H_{B},$ (4)
in which $N_{A}$ is some “number” operator in $A$ with eigenvalues $\{0,1,\ldots,d_{A}-1\}$ and eigenstates $\lvert n\rangle_{A}$. Interactions of this type can be engineered, for example, in quantum dots [39] and Rydberg atoms [22]. We will take $A$ to be initialized to an equal superposition of $N_{A}$-eigenstates, as is conventional in such protocols:
$\lvert\chi\rangle_{A}=\frac{1}{\sqrt{d_{A}}}\sum_{n=0}^{d_{A}-1}\lvert n\rangle_{A}.$ (5)
We will also assume that $B$ is initialized to an unspecified initial state $\lvert\phi\rangle_{B}$ (owing to Bob’s limited control), whose properties determine certain advantages of our protocol.
The state of the combined system at time $t_{0}$ is
$\lvert\Psi(t_{0})\rangle=e^{-iHt_{0}}\lvert\chi\rangle_{A}\otimes\lvert\phi\rangle_{B}.$ (6)
Defining the unitary operator $U_{t_{0}}=\exp(-iH_{B}t_{0})$, we see that this time-dependent state can be expressed as:
$\lvert\Psi(t_{0})\rangle=\frac{1}{\sqrt{d_{A}}}\sum_{n=0}^{d_{A}-1}\lvert n\rangle_{A}\otimes U_{t_{0}}^{n}\lvert\phi\rangle_{B}.$ (7)
From the arguments pertaining to Eq. (3), it is clear that the closeness of $\lvert\Psi(t_{0})\rangle$ to an EPR state is decided by the orthonormality of the $d_{A}$-element set of Krylov vectors [57, 58, 59] generated by successive actions of $U_{t_{0}}$ on $\lvert\phi\rangle_{B}$:
$\mathcal{K}_{d_{A}}(U_{t_{0}},\lvert\phi\rangle_{B})=\left\{\lvert\phi\rangle_{B},U_{t_{0}}\lvert\phi\rangle_{B},\ldots,U_{t_{0}}^{d_{A}-1}\lvert\phi\rangle_{B}\right\}.$ (8)
These vectors track [58] the dynamics of $\lvert\phi\rangle_{B}$, and span the subspace $\mathcal{H}_{2}\subseteq\mathcal{H}_{B}$ in which Bob can encode information from Alice.
At a conceptual level, the orthonormality of the Krylov set is connected to spectral statistics as follows. In Ref. [37], it is shown (adapted to our present language) that the smallness of spectral fluctuations, such as for random matrices, allows the existence of an orthonormal basis in $\mathcal{H}_{B}$ such that a $d_{B}$-element Krylov set of each basis element has a strong overlap with the original basis, which provides a quantum notion of dynamical ergodicity (in the sense of dynamics exploring a complete orthonormal basis). While it is difficult to make a direct quantitative connection of our present protocol with this notion, the conceptual mechanism at play here is similar: we will see that suppressed spectral fluctuations, both of the energy levels and of the initial state $\lvert\phi\rangle_{B}$, lead to an increased orthonormality of the Krylov set.
We call the latter phenomenon “Krylov vector ergodicity”. As an aside, we note its potential to bridge the notions of dynamical ergodicity [37] and Krylov state complexity [60, 61] (delocalization in a Krylov set related to actions of $H_{B}$ rather than $U_{t_{0}}$), whose correspondence has been conjectured in Ref. [38] based on theoretical quantum gravity observations.
Ergodic entanglement generation— We will measure entanglement between $A$ and $B$, and therefore the orthonormality of the Krylov set in Eq. (8), by means of the purity [47] (which is the exponential of the negative of the $2$nd Rényi entanglement entropy [62, 63]):
$\mathcal{P}(t)=\operatorname{Tr}_{B}\left[\left\{\vphantom{\sum}\operatorname{Tr}_{A}\left(\lvert\Psi(t)\rangle_{AB}\langle\Psi(t)\rvert\right)\right\}^{2}\right].$ (9)
In all cases, we have $d_{A}^{-1}\leq\mathcal{P}(t)\leq 1$, with the lower limit $d_{A}^{-1}$ corresponding to maximal entanglement as for EPR states. For the state in Eq. (7), it follows that
$\mathcal{P}(t_{0})=\frac{1}{d_{A}}+\frac{2}{d_{A}}\sum_{\tau=1}^{d_{A}-1}\left(1-\frac{\tau}{d_{A}}\right)p_{\phi}(\tau t_{0}),$ (10)
where the deviation from maximal entanglement is directly given by the dynamics of the return probability $p_{\phi}(t)$ of the initial state:
$p_{\phi}(t)$ $\equiv\left\lvert\langle\phi\rvert e^{-iH_{B}t}\lvert\phi\rangle_{B}\right\rvert^{2}$
$=\sum_{n,m=0}^{d_{B}-1}\lvert\phi(E_{n})\rvert^{2}\lvert\phi(E_{m})\rvert^{2}e^{-i(E_{n}-E_{m})t}.$ (11)
Here, $\lvert E_{n}\rangle_{B}$ refers to the energy eigenstates of $H_{B}$ with respective eigenvalue $E_{n}$, and $\phi(E_{n})=\langle E_{n}|\phi\rangle_{B}$ are the components of the initial state in this basis. The essential intuition behind Eq. (10) is that a larger overlap with the initial state over time hinders the ability of dynamics to effectively explore the rest of the Hilbert space to generate more orthonormal states [37], therefore increasing purity and reducing entanglement.
The quantity $p_{\phi}(t)$ is sensitive to both the fluctuations of the energy levels $E_{n}$, and the fluctuations of the state components $\phi(E_{n})$, in a manner that has been studied in several works, e.g. [64, 65]. The details of this behavior are reviewed in App. B; here, we note the relevant qualitative trends. After a time $t>t_{\text{ramp}}$, determined by the energy scale of fluctuations, we can write $p_{\phi}(t)$ as a sum of contributions from spectral fluctuations and initial state fluctuations:
$p_{\phi}(t>t_{\text{ramp}})\sim f_{E}(t)+f_{\phi}(t).$ (12)
The contribution from spectral statistics is, for $t$ not too long:
$f_{E}(t)\sim\begin{dcases}O(t/d_{B}^{2}),&\ \text{ergodic $E_{n}$},\\ 1/d_{B},&\ \text{Poisson $E_{n}$},\end{dcases}$ (13)
reflecting the smaller spectral fluctuations for ergodic systems (typically, with random matrix statistics). The contribution from the initial state depends on whether $\phi(E)$ varies smoothly as a function of energy $E$. This is the case for special initial states, such as discrete Fourier transforms of the energy eigenstates, or coherent Gibbs states [66]. More generic initial states in complex systems, however, fluctuate randomly. We have:
$f_{\phi}(t)\sim\begin{dcases}0,&\ \text{smooth }\phi(E),\\ 1/d_{B},&\ \text{generic }\phi(E).\end{dcases}$ (14)
The combination of random matrix statistics (or smaller spectral fluctuations) and a smooth $\phi(E)$ corresponds to a highly ergodic Krylov set. In consequence, from Eq. (10), we have for the purity at the (not too long) time $t_{0}>t_{\text{ramp}}$:
$\mathcal{P}(t_{0})=\frac{1}{d_{A}}+\begin{dcases}O\left(d_{A}/d_{B}^{2}\right),&\ \text{ergodic $E_{n}$ \& smooth $\phi$},\\ 1/d_{B},&\ \left\{\begin{subarray}{c}\text{ergodic $E_{n}$ \& generic $\phi$},\\ \text{Poisson $E_{n}$ \& smooth $\phi$},\end{subarray}\right.\\ 2/d_{B},&\ \text{Poisson $E_{n}$ \& generic $\phi$}.\end{dcases}$ (15)
It is worth comparing these results to entanglement generation via scrambling dynamics. A typical “scrambling” Hamiltonian $H_{s}$ with local interactions between $A$ and $B$ generally thermalizes states to a finite temperature $\beta^{-1}$ at all sufficiently long times $t>t_{s}$ (up to quantum recurrences [67, 68]), incapable of generating maximal entanglement. Quantitatively, Ref. [69] suggests that for generic finite temperature local Hamiltonian dynamics, one may have
$\mathcal{P}_{\beta}(t>t_{s})\sim d_{A}^{-1+c\beta^{2}},$ (16)
where $c>0$ is a nonvanishing $O(1)$ parameter that depends on $d_{A}$ and $d_{B}$, which is not maximal entanglement even to leading order for $\beta>0$. Infinite temperature scrambling to $\beta=0$, such as for strongly nonlocal Hamiltonians [70, 71, 44] or random quantum circuits [72], can generate maximal entanglement up to subleading corrections:
$\mathcal{P}_{\beta=0}(t>t_{s})=\frac{1}{d_{A}}+\frac{1}{d_{B}},$ (17)
as estimated for ideal random systems [73, 74, 75, 72, 23] (neglecting higher order terms for $d_{A}\ll d_{B}$, including subleading spectral corrections [74, 75, 65]). Larger deviations from minimal purity may occur for physical nonlocal Hamiltonians [71, 70]. From Eq. (15), even with a generic random wavefunction $\phi(E)$ as may be prepared in typical many-body systems, a random matrix spectrum produces the same entanglement as Eq. (17) in our protocol. Thus, in generic cases, ergodic entanglement generation via Eq. (4) is at least as good as an infinite temperature scrambler even if $H_{B}$ is not strongly interacting within Bob’s system, and with specific “smooth” initial states, is considerably closer to a maximally entangled state as per the top row of Eq. (15).
Operator transfer capacity— Having identified an enhancement in entanglement generation from ergodic dynamics, it is worth asking to what extent this enhancement actually matters in quantum information processing. We will find a notable advantage for operator transfer as in Eq. (2), implying corresponding gains in tasks such as quantum teleportation (relative to scrambling dynamics).
To understand these gains, it is useful to write our entangled state as the EPR state with a pre-loaded operator:
$\lvert\Psi(t)\rangle=R_{B}(t)\lvert\mathrm{EPR}\rangle,$ (18)
where $R_{B}$ is a general linear transformation in $\mathcal{H}_{B}$, transforming some orthonormal basis to the states $\lvert\psi_{k}\rangle_{2}$ in Eq. (3), such as the Krylov set in Eq. (8). For example, this transformation determines the purity via $\mathcal{P}(t)=\operatorname{Tr}_{B}[(R_{B}(t)R_{B}^{\dagger}(t))^{2}]/d_{A}^{2}$. We can now identify a generalized operator transfer property:
$O_{A}\lvert\Psi(t)\rangle=R_{B}(t)O_{B}^{T}\lvert\mathrm{EPR}\rangle,$ (19)
where $O_{A}$ in Alice’s system, acting on our entangled state, behaves as if it is mapped to $R_{B}(t)O_{B}^{T}$ in Bob’s system by an ideal EPR state. For faithful transfer, we demand that the trace inner products of operators are approximately preserved; for any two operators $O_{A}$, $P_{A}$ with Alice, we require:
$\operatorname{Tr}_{B}\left[\left\{R_{B}(t)O_{B}^{T}\right\}\left\{R_{B}(t)P_{B}^{T}\right\}^{\dagger}\right]\simeq\operatorname{Tr}_{A}\left[O_{A}P_{A}^{\dagger}\right].$ (20)
This allows teleportation via a $\langle\Psi(t)\rvert$-measurement of $B$ with a third system [46, 47, 48, 49]; related ideas will be discussed in upcoming work [76]. In App. C, we formulate this criterion more precisely to allow for a relative error $10\leq\epsilon<$ 1 in this inner product. For successful operator transfer with an error of at most $\epsilon$, we find that the purity must satisfy:
$\mathcal{P}(t)\leq\frac{1}{d_{A}}+\frac{\epsilon^{2}}{d_{A}^{\gamma}(\ln d_{A})^{\kappa}},$ (21)
With $\kappa=0$, for $\gamma=1$, this gives a strict necessary condition for successful operator transfer  [^1] The maximal scrambling criterion used in [44, 45] is in turn a necessary condition for this to hold.; for $\gamma=2$ we get a rigorous sufficient condition. We also get a “typical” condition, perhaps necessary and sufficient in several generic cases, when $\gamma=1$, $\kappa\sim 1$. A more specific analysis requires incorporating higher order interference effects than what is visible to $p_{\phi}(t)$ [see Eqs. (29) and (34)], which appears to be a considerable challenge, and may provide an interesting direction for future work.
Let us now use Eq. (21) to contrast ergodic entanglement generation with scramblers. By Eq. (16), typical local Hamiltonian scramblers do not generate sufficient entanglement for any operator transfer as their purity saturates at $\gamma\sim 1-c\beta^{2}<1$. But a comparable Hamiltonian realized in Bob’s system via Eq. (4) is capable of operator transfer due to ergodicity, as are infinite temperature $\beta=0$ scramblers. For these systems, the subleading terms in Eqs. (15) and (17) set a direct lower limit on the dimension $d_{B}$ of Bob’s system required to faithfully encode information from Alice, where we regard a smaller lower limit on $d_{B}$ as a diagnostic of increased operator transfer capacity:
$d_{B}\gtrsim\begin{dcases}[d_{A}^{(1+\gamma)/2}(\ln d_{A})^{\kappa/2}]/\epsilon,&\ \text{ergodic $E_{n}$ \& smooth $\phi$},\\ [\varphi d_{A}^{\gamma}(\ln d_{A})^{\kappa}]/\epsilon^{2},&\ \text{generic $\phi$, OR $\beta=0$}.\end{dcases}$ (22)
In the second line, $\varphi=1$ for a random matrix spectrum or ideal $\beta=0$ scrambler, while $\varphi=2$ for Poisson statistics, indicating a mildly improved capacity for the former. However, the first line is parametrically smaller for $\gamma\geq 1$ as in Eq. (21) (given $\epsilon<1$). Consequently, with full ergodicity (i.e., an ergodic spectrum and a smooth initial state) and the sufficient condition in Eq. (21), we can have $d_{B}\sim d_{A}^{3/2}/\epsilon$, but otherwise require $d_{B}\gtrsim(d_{A}/\epsilon)^{2}$, indicating that ergodicity can provide up to a power-law gain. More generically, with the “typical” parameters of Eq. (21), full ergodicity can faithfully encode Alice’s information into a logarithmically smaller system of Bob, $d_{B}\sim d_{A}\sqrt{\ln d_{A}}/\epsilon$, than with a generic infinite temperature scrambler or Poisson statistics for which $d_{B}\gtrsim d_{A}\ln d_{A}/\epsilon^{2}$. Even with generic initial states, a factor of $2$ advantage over Poisson statistics is retained. These cases illustrate a clear advantage of ergodicity in operator transfer capacity (via the smallness of Bob’s system), which carries over into protocols that rely on the EPR state such as quantum teleportation.
Discussion— We have shown that ergodicity can be harnessed, even in the standard context of generating EPR states, to generate parametrically higher entanglement than scrambling dynamics for tasks such as quantum teleportation. This suggests a practical motivation for studying the spectral statistics of large-but-not-thermodynamically-large systems, to realize a simple goal of maximizing operator transfer capacity in highly entangled states. We regard this possibility as surprising to the extent that physical applications of even classical ergodic dynamics are scarce, outside formal theoretical insights [78].
Most importantly, our approach emphasizes a shift in viewpoint, from considering spectral statistics as a qualitative “defining” notion for “quantum chaos”, whose implications remain rather vague, to a resource for quantum information processing with a demonstrable theoretical advantage. It is conceivable that other operational applications of spectral statistics exist, perhaps more significant than outlined here, that deserve to be explored; see also [79] for the significance of operational notions in quantum information.
In conclusion, it is important to emphasize a qualifier. Our approach should be regarded as a theoretical proof-of-principle for the practical utility of spectral statistics, but its potential for complete experimental realization remains to be seen. The primary challenge is the complexity of preparing a “smooth” initial state that shows maximally ergodic dynamics, corresponding to the top row of Eq. (14). Nevertheless, even without the ability to easily prepare such a state, which diminishes its advantage over (ideal) infinite temperature scramblers, our protocol continues to have the potential for present-day use cases. A notable feature is its ability to generate enhanced entanglement comparable to infinite temperature scramblers even with generic initial states in an ergodic system, while continuing to use e.g. a local finite-temperature scrambling Hamiltonian for Bob’s system in Eq. (4) that may otherwise be incapable of generating such entanglement.
Acknowledgements.
We thank Andrew Lucas for useful discussions. We acknowledge related collaborations with Chris Akers, Edwin Chaparro, Muhammad Miskeen Khan, Andrew Lucas, and Ana Maria Rey. This work was supported by the Heising-Simons Foundation under Grant 2024-4848.
Appendix A Multidimensional Krylov vector ergodicity
The most general nondemolition interaction between $A$ and $B$, of which Eq. (4) is a special case, is a Hamiltonian that couples conserved charges between the two systems:
$H=H_{A}\otimes\mathds{1}_{B}+\mathds{1}_{A}\otimes H_{B}+\sum_{k}q_{Ak}\otimes Q_{Bk}.$ (23)
Here, $H_{A}$ and $H_{B}$ represent the independent dynamics of $A$ and $B$, while $q_{Ak}$ and $Q_{Bk}$ are mutually commuting conserved charges: $[q_{Ak},H_{A}]=[q_{Ak},q_{Aj}]=0$, and $[Q_{Bk},H_{B}]=[Q_{Bk},Q_{Bj}]=0$. Eq. (23) may be easier to implement in a wider range of systems [39, 80] than Eq. (4).
Let $\lvert n\rangle_{A}$ and $\lvert E_{m}\rangle_{B}$ be joint eigenbases of these operators in $A$ and $B$, with $q_{Ak}$ eigenvalues $q_{Ak,n}$ and $Q_{Bk}$ eigenvalues $Q_{Bk,m}$. Then for the same dynamics as Eq. (6), we obtain
$\lvert\Psi(t_{0})\rangle=\frac{e^{-iH_{0}t_{0}}}{\sqrt{d_{A}}}\sum_{n=0}^{d_{A}-1}\lvert n\rangle_{A}\ \otimes\prod_{k}e^{-i(q_{Ak,n}t_{0})Q_{Bk}}\lvert\phi\rangle_{B},$ (24)
where $H_{0}=H_{A}\otimes\mathds{1}_{B}+\mathds{1}_{A}\otimes H_{B}$. The relevant $d_{A}$-element Krylov set is now determined by the multidimensional group of symmetry transformations $U_{t_{0},k}=\exp(-iQ_{Bk}t_{0})$ generated by the different conserved charges, over $d_{A}$ different sets of multidimensional “angles” $q_{Ak,n}$ determining the duration of each generating transformation:
$\mathcal{K}_{q_{Aj}}\left(U_{t_{0},k},\lvert\phi\rangle_{B}\right)=\left\{\lvert\mathcal{K}_{n}\rangle=\prod_{k}U_{t_{0},k}^{q_{Ak,n}}\lvert\phi\rangle_{B},\ \forall n\in\mathbb{Z}_{d_{A}}\right\}.$ (25)
The orthonormality of this basis is sensitive to the spectral statistics of each charge $Q_{Bk}$. For sufficiently complex charges, the $Q_{Bk}$ can have nontrivial spectral statistics. For the setting of Refs. [39, 80] where $B$ is a quantum dot, the charge $Q_{B1}$ is a randomly weighted sum of local spins, and may generically have Poisson statistics. If any $Q_{Bk}$ corresponds to a sufficiently complex operator in $B$, then it may have random matrix statistics.
Recall that, by definition, the Krylov set spans a subspace $\mathcal{H}_{2}\subseteq\mathcal{H}_{B}$ of dimension $d_{2}\leq d_{A}$, as we have $d_{A}$ Krylov vectors. We now extend $\mathcal{H}_{2}$ by introducing auxiliary dimensions if necessary to make $d_{2}=d_{A}$. If $\lvert j\rangle_{2}$ is an orthonormal basis for this extended $\mathcal{H}_{2}$ space, we can write the Krylov set as a general linear transformation of this orthonormal basis:
$\lvert\mathcal{K}_{j}\rangle_{2}=R_{2}\lvert j\rangle_{2}.$ (26)
In the ideal case where the Krylov set is orthonormal, $R_{2}$ is a unitary operator satisfying $R_{2}R_{2}^{\dagger}=R_{2}^{\dagger}R_{2}=\mathds{1}_{2}$. More generally, we can measure “Krylov vector ergodicity” via the closeness of $R_{2}$ to a unitary transformation, e.g. the smallness of
$\eta_{2}\equiv\frac{1}{d_{2}}\operatorname{Tr}_{2}\left[(R_{2}R_{2}^{\dagger})^{2}\right]-1,$ (27)
with $\eta_{2}=0$ for a perfect unitary transformation (which generates ideal EPR states) and $\eta_{2}>0$ otherwise. This is because $R_{2}R_{2}^{\dagger}$ and $R_{2}^{\dagger}R_{2}$ are positive operators (with nonnegative eigenvalues) and $\operatorname{Tr}_{2}[R_{2}R_{2}^{\dagger}]=d_{2}$ is fixed by $\langle\mathcal{K}_{j}|\mathcal{K}_{j}\rangle=1$, which follows from Eq. (25); therefore, $\eta_{2}=0$ implies that $R_{2}R_{2}^{\dagger}=R_{2}^{\dagger}R_{2}=\mathds{1}_{2}$. We emphasize that, unlike dynamical ergodicity in Ref. [37] which is either present or absent in a system (which is useful to make quantitative connections to random matrix statistics), it is more convenient to regard Krylov vector ergodicity as a “continuous” measure whose degree (of absence) is given by $\eta_{2}$. An explicit calculation of $\eta_{2}$ verifies its connection to multidimensional return probabilities of $\lvert\phi\rangle_{B}$ under the symmetry transformations $U_{t_{0},k}$, generalizing Eq. (10).
Finally, if we evaluate the purity $\mathcal{P}(t_{0})$ of the state $\lvert\Psi(t_{0})\rangle$ according to Eq. (9), we get
$\mathcal{P}(t_{0})=\frac{1}{d_{A}^{2}}\operatorname{Tr}_{2}\left[(R_{2}R_{2}^{\dagger})^{2}\right]=\frac{1+\eta_{2}}{d_{A}},$ (28)
verifying that the deviation from minimal purity $1/d_{A}$ (or maximal entanglement) directly measures Krylov vector ergodicity in the above sense. More generally, we can work with “higher purities” (related to the $\alpha$-Rényi entropy [62], with $\alpha=2$ giving the conventional purity),
$\mathcal{P}_{\alpha}(t_{0})\equiv\operatorname{Tr}_{A}\left[\rho^{\alpha}_{A}(t_{0})\right]=\frac{1}{d_{A}^{\alpha}}\operatorname{Tr}_{2}\left[(R_{2}R_{2}^{\dagger})^{\alpha}\right],$ (29)
where $\rho_{A}(t_{0})=\operatorname{Tr}_{B}[\lvert\Psi(t_{0})\rangle\langle\Psi(t_{0})\rvert]$, which are given by e.g. more complicated functions of the amplitudes $\langle\phi\rvert U_{t_{0}}^{\tau_{j}}\lvert\phi\rangle_{B}$ than $p_{\phi}(t)$, involving multiple times $\tau_{j}$ with a cyclicity condition $\sum_{j}\tau_{j}=0$ and some associated phase interference effects for the case of Eq. (4); we will therefore leave a detailed study of these quantities (for $\alpha\neq 2$) for future work.
Appendix B Spectral fluctuations in return probabilities
Now, we will review some quantitative details of the behavior of $p_{\phi}(t)$ in Eq. (11), and some miscellaneous issues. As there are much better resources discussing this behavior with more scope than this short Appendix, we will primarily just point to the literature here with brief accompanying details.
Review— At early times, $p_{\phi}(t)$ decays smoothly from $1$ at $t=0$, in a way that depends on the coarse-grained profile of $\phi(E)$. When $p_{\phi}(t)\sim O(1/d_{B})$ after a time $t>t_{f}$, it fluctuates erratically [81] depending on the $E_{n}$ and $\phi(E_{n})$. This is usually not analytically tractable, but reasonable estimates have been made in the literature [64, 65, 82] assuming that these fluctuations are sufficiently random.
In most systems, to a good approximation, these are often independent, and add up independently as in Eq. (12). As Eq. (10) sums over $p_{\phi}(t)$ at different times, it is reasonable to approximate these (non-negative) fluctuations by their average value, as we have done above. The contribution $f_{E}(t)$ may be isolated (up to fluctuations) by setting $\phi(E_{n})=1/d_{B}$, for which $f_{\phi}(t)=0$; for this choice, $p_{\phi}(t)$ exactly corresponds to the state-independent “spectral form factor”, whose phenomenology has been explored in detail in e.g. [1, 2, 12, 8, 83, 9, 10, 11, 13, 14, 22, 23, 24, 37, 44, 45]. After a system-specific [12] time $t_{\text{ramp}}>t_{f}$ (called the “ramp” or “Thouless” time), which may be [84] as small as $O(1)$ and smaller or larger than the scrambling time $t_{s}$, $f_{E}(t)\sim t/d_{B}^{2}$ (the “ramp”) for a random matrix spectrum up to the Heisenberg time $t_{{\rm H}}\sim 2\pi\mathcal{N}(E)$ determined by the typically $O(d_{B})$ density of states $\mathcal{N}(E)$. Our “ergodic $E_{n}$” expression in Eq. (13) assumes that $t\ll t_{{\rm H}}$, including for $t=t_{0}d_{A}$, quantifying which times we regard as “not too long”. Poisson statistics corresponds to a constant $f_{E}(t)\sim 1/d_{B}$ shortly after $t_{f}$, as does random matrix statistics after $t\sim t_{{\rm H}}$. For simplicity, we neglect other cases of intermediate statistics.
The wavefunction contribution $f_{\phi}(t)$ has been considered in e.g. [64, 65, 82], and is generally smaller for more delocalized wavefunctions in $E$. Eq. (14) corresponds to fully delocalized wavefunctions with maximal randomness, as for a mathematically (Haar) random state, in the second row. Any reduction in randomness would decrease this offset. The first row with $f_{\phi}(t)\sim 0$ includes $\phi(E_{n})=1/d_{B}$, but also other smooth wavepackets such as $\phi(E)\sim\exp(-(E-E_{0})^{2}/4\sigma^{2})$ or $\phi(E)\sim\exp(-\beta E/2)$; the absence of fluctuations leads to negligible wavefunction contributions a while after $t_{f}$.
Misc. details— To generate maximal entanglement, we would ideally like the range of times $[t_{0},t_{0}d_{A}]$ to be as close as possible to $t_{\text{ramp}}$ for minimal $f_{E}(t)$. A spectrum or $\phi(E)$ with sharp features can obscure the ramp time due to slowly decaying oscillations [12, 44], so a smooth wavepacket is also our best bet for the earliest detectable $t_{\text{ramp}}$. Discrete-time systems with a bounded spectrum such as quantum circuits are generically subject to such oscillations [44] unless one can fine-tune $t_{0}$ and $N_{A}$, so a continuous time Hamiltonian may tolerate errors in $t_{0}$ or $N_{A}$ better.
To avoid state preparation issues, we note that if we initialize $d_{B}$ to a maximally mixed state $\rho_{B}=\mathds{1}_{B}/d_{B}$ (which is easily prepared by maximal scrambling with a larger system $\mathcal{H}_{R}$ that may be ignored after state preparation [54]) instead of a pure state $\lvert\phi\rangle_{B}$, then $p_{\phi}(t)$ in Eq. (28) is replaced by the exact spectral form factor $K(t)=\lvert\operatorname{Tr}_{B}(e^{-iH_{B}t})/d_{B}\rvert^{2}$. The resulting purity is $\mathcal{P}(t_{0})=1/d_{A}+O(d_{A}/d_{B}^{2})$; however, an infinite temperature scrambler generates $\mathcal{P}(t>t_{s})=1/d_{A}+O(1/d_{B}^{2})$, which is more efficient. Nevertheless, this scenario may still be useful to formally generate near-minimal purity in $A$ by leveraging the dynamical ergodicity of finite temperature e.g. local Hamiltonians $H_{B}$ in Eq. (4) compared to Eq. (16), without state preparation difficulties, but also without any operator transfer capacity from $A$ to $B$ due to the mixed initial state.
Appendix C Details of operator transfer capacity
Here, we quantify the criterion for faithful operator encoding in Eq. (20). By the Cauchy-Schwarz inequality,
$\left\lvert\operatorname{Tr}_{A}[O_{A}P_{A}^{\dagger}]\right|\leq\sqrt{\operatorname{Tr}_{A}[O_{A}O_{A}^{\dagger}]\operatorname{Tr}_{A}[P_{A}P_{A}^{\dagger}]}.$ (30)
We will let the right hand side set the scale relative to which we measure the error $\epsilon$ in the inner products of encoded operators $O_{A}\to R_{2}O_{2}^{T}$, in the notation of App. A. We consider the entangled state $\lvert\Psi(t_{0})\rangle$ to be able to successfully encode operators with an error of $\epsilon$ or less if $\Delta_{2}(O_{A},P_{A})\leq\epsilon$, where
$\Delta_{2}(O_{A},P_{A})\equiv\frac{\left\lvert\operatorname{Tr}_{2}\left[\left\{R_{2}O_{2}^{T}\right\}\left\{R_{2}P_{2}^{T}\right\}^{\dagger}\right]-\operatorname{Tr}_{A}[O_{A}P_{A}^{\dagger}]\right\rvert}{\sqrt{\operatorname{Tr}_{A}[O_{A}O_{A}^{\dagger}]\operatorname{Tr}_{A}[P_{A}P_{A}^{\dagger}]}}.$ (31)
As $\mathcal{H}_{2}$ is isomorphic to $\mathcal{H}_{A}$ by construction, it follows that (with some implicit transposes and cycling of operators):
$\Delta_{2}(O_{A},P_{A})=\frac{\left\lvert\operatorname{Tr}_{2}\left[(R_{2}^{\dagger}R_{2}-\mathds{1}_{2})O_{2}^{T}P_{2}^{T\dagger}\right]\right\rvert}{\sqrt{\operatorname{Tr}_{A}[O_{A}O_{A}^{\dagger}]\operatorname{Tr}_{A}[P_{A}P_{A}^{\dagger}]}}.$ (32)
Let us first take $O_{A}=P_{A}$. Then, $(O_{2}^{\dagger}O_{2})^{T}$ is a positive operator with nonnegative eigenvalues $\omega_{j}\geq 0$ and orthonormal eigenstates $\lvert\omega_{j}\rangle$. The above expression becomes:
$\Delta_{2}(O_{A},O_{A})=\frac{\sum_{j}\omega_{j}\left\lvert\langle\omega_{j}\rvert(R_{2}^{\dagger}R_{2}-\mathds{1}_{2})\lvert\omega_{j}\rangle\right\rvert}{\sum_{j}\omega_{j}}.$ (33)
Consequently, a necessary and sufficient condition for $\Delta_{2}(O_{A},O_{A})\leq\epsilon$ for any $O_{A}$ is that
$\max_{\lvert\psi\rangle\in\mathcal{H}_{2}:\ \langle\psi|\psi\rangle=1}\ \left\lvert\langle\psi\rvert(R_{2}^{\dagger}R_{2}-\mathds{1}_{2})\lvert\psi\rangle\right\rvert\leq\epsilon.$ (34)
This measure can, in principle, be accessed through the higher purities $\mathcal{P}_{\alpha}(t_{0})$ in Eq. (29), for example $\lim_{\alpha\to\infty}\mathcal{P}_{\alpha}^{1/\alpha}(t_{0})$. However, as the behavior of these quantities is complicated to evaluate, we will attempt to relate Eq. (34) to $\mathcal{P}(t_{0})$. Let $r_{k}\geq 0$ denote the eigenvalues of $R_{2}^{\dagger}R_{2}$, whose mean is $1$ as $\operatorname{Tr}_{2}[R_{2}R_{2}^{\dagger}]/d_{2}=1$; the above condition enforces ${\lvert r_{k}-1\rvert}\leq\epsilon$. Then, $\eta_{2}=d_{2}^{-1}\sum_{k}(r_{k}-1)^{2}$ measures the variance of the $r_{k}$ around $1$, and the Bhatia-Davis inequality [85] gives
$\eta_{2}\leq(r_{\max}-1)(1-r_{\min})\implies\eta_{2}\leq\epsilon^{2},$ (35)
where $r_{\max}$ and $r_{\min}$ are the maximum and minimum of the $r_{k}$, as a necessary condition for Eq. (34) to be valid, implying Eq. (21) with $\gamma=1$, $\kappa=0$. For a sufficient condition based on $\eta_{2}$, the best we can do is impose that $\eta_{2}\leq\epsilon^{2}/d_{2}$, corresponding to Eq. (21) with $\gamma=2$ and $\kappa=0$. If the $r_{k}$ follow a Gaussian distribution, which is likely to be generic, then we expect $(r_{\max}-1)$ and $(1-r_{\min})$ to be [86, 87, 88] near $\sqrt{2\eta_{2}\ln d_{2}}$, corresponding to Eq. (21) with $\gamma=1$ and $\kappa\sim 1$.
For $O_{A}\neq P_{A}$, expanding $R_{2}^{\dagger}R_{2}$ in its orthonormal eigenbasis $\lvert r_{k}\rangle$ and using Eq. (34) with the triangle inequality gives:
$\Delta_{2}(O_{A},P_{A})\leq\epsilon\frac{\max_{k}\left\lvert\langle r_{k}\rvert O_{A}P_{A}^{\dagger}\lvert r_{k}\rangle\right\rvert}{\sqrt{\operatorname{Tr}_{A}[O_{A}O_{A}^{\dagger}]\operatorname{Tr}_{A}[P_{A}P_{A}^{\dagger}]}}$ (36)
As $\lvert\langle\psi\rvert OP\lvert\psi\rangle\rvert\leq\sqrt{\langle\psi\rvert OO^{\dagger}\lvert\psi\rangle\langle\psi\rvert PP^{\dagger}\lvert\psi\rangle}$ by the Cauchy-Schwarz inequality, with $OO^{\dagger}$ and $PP^{\dagger}$ being positive operators, we get that $\Delta_{2}(O_{A},P_{A})\leq\epsilon$ if Eq. (34) is satisfied.